OpenStudy (hiroromon):

Can someone teach me how to verify that this equation is an identity step by step? cosA - sinA = (cos2A)/(cosA + sinA)

OpenStudy (raden):

that's easy, use the identity : cos2A = cos^2 A - sin^2 A

OpenStudy (raden):

see your right side becomes, (cos2A)/(cosA + sinA) = (cos^2 A - sin^2 A)/(cosA + sinA)

OpenStudy (raden):

then factor out the numerator by using the difference square formula : a^2 - b^2 = (a+b)(a-b)

OpenStudy (hiroromon):

ohhhh

OpenStudy (hiroromon):

i get it, thank you :)

OpenStudy (raden):

so, it can be = (cos A - sinA)(cos A + sinA)/(cosA + sinA) cancel out the similar term ;)

OpenStudy (raden):

you're welcome